OpenStudy (anonymous):

Factorise the expression, 14w^2+23w+3 and hence solve the equation 14w^2+23w+3=0.

OpenStudy (aroub):

(7w+1) (2x+3) so w= -3/2 and w= -1/7

jhonyy9 (jhonyy9):

there is w and not x inside second paranthese 2w+3

OpenStudy (aroub):

oops! sorry

jhonyy9 (jhonyy9):

no problem just for who have asked

OpenStudy (anonymous):

i'm still confused because the equation doesn't include anything that gives the answer 23w :S

jhonyy9 (jhonyy9):

make the calcule newly 21w+2w=23w